### v1

# MAGE: Model-Level Graph Neural Networks Explanations via Motif-based Graph Generation

Abstract
Graph Neural Networks (GNNs) have shown remarkable success in molecular tasks, yet their interpretability remains challenging. Traditional model-level explanation methods like XGNN and GNNInterpreter often fail to identify valid substructures like rings, leading to questionable interpretability. This limitation stems from XGNN’s atom-by-atom approach and GNNInterpreter’s reliance on average graph embeddings, which overlook the essential structural elements crucial for molecules. To address these gaps, we introduce an innovative Motif-bAsed GNN Explainer (MAGE) that uses motifs as fundamental units for generating explanations. Our approach begins with extracting potential motifs through a motif decomposition technique. Then, we utilize an attention-based learning method to identify class-specific motifs. Finally, we employ a motif-based graph generator for each class to create molecular graph explanations based on these class-specific motifs. This novel method not only incorporates critical substructures into the explanations but also guarantees their validity, yielding results that are human-understandable. Our proposed method’s effectiveness is demonstrated through quantitative and qualitative assessments conducted on six real-world molecular datasets.
1 Introduction
Graph Neural Networks (GNNs) [18, 39, 12, 28, 33, 38] have become increasingly popular tools for modeling data in the molecule field. As an effective method for learning representations from molecule data, GNNs have attained state-of-the-art results in tasks like molecular representation learning [11, 43, 10, 48], and molecule generation [6, 20]. Despite their growing popularity, questions arise about the trustworthiness and decision-making processes of GNNs.
Explaining GNNs has become a major area of interest in recent years and existing methods can be broadly categorized into instance-level explanations and model-level explanations. Instance-level explanations [42, 4, 26, 13, 34, 29, 22, 49, 30, 36, 47, 35, 44] aim to pinpoint specific nodes, edges, or subgraphs crucial for a GNN model’s predictions on one data instance. Model-level explanations [46, 37] seek to demystify the overall behavior of the GNN model by identifying patterns that generally lead to certain predictions. While instance-level methods offer detailed insights, they require extensive analysis across numerous examples to be deemed reliable. Conversely, model-level methods provide more high-level explanations that require less intensive human oversight.
Model-level explanation methods have two main categories: concept-based methods and generation-based methods. Concept-based explanation methods focus on identifying higher-level concepts that significantly influence the model’s predictions [3, 40]. Also, these methods establish rules to illustrate how these concepts are interconnected, like using logical formulas [3]. Generation-based models try to learn a generative model that can generate synthetic graphs that are optimized to maximize the behavior of the target GNN [46, 37]. Instead of defining the relationships between important concepts with formulas, generation-based models can generate novel graph structures that are optimized for specific properties. This capability is particularly crucial in the field of molecular learning.
Figure 1: Validity of explanations generated by XGNN and GNNInterpreter on Mutagenicity dataset
Despite the notable benefits of generation-based model-level explanation methods for GNNs, they have received relatively low attention. Presently, the available generation-based model-level explainers for GNNs are somewhat limited in their application, especially in molecular graphs. For example, XGNN [46] explains models by building an explanation atom-by-atom and sets a maximum degree for each atom to maintain the explanation’s validity. On the other hand, GNNInterpreter [37] adopts a more flexible approach by learning a generative explanation graph distribution. However, it ensures the validity of explanations by maximizing the similarity between the explanation graph embedding and the average embedding of all graphs, which is not particularly effective for molecular graphs. These existing methods often fail to consider the unique aspects of molecular structures adequately. This oversight makes it challenging for these methods to include crucial substructures in their explanation graphs. It compromises the validity and reliability of the model-level explanations they provide for molecular graphs. In Figure 1, the explanations produced by both XGNN and GNNInterpreter exhibit low validity, making them ineffective for further analysis. Consequently, there’s a clear need for more advanced and specialized model-level explanation methods that can effectively handle the unique characteristics of molecular graphs and provide meaningful insights into how GNNs interpret and process molecular data.
This paper introduces a motif-based approach for model-level explanations, MAGE, specifically tailored for GNNs in molecular representation learning tasks. The method begins by identifying all possible motifs from the molecular data. Following this, our method focuses on pinpointing particularly significant motifs for a specific class. This is achieved through a novel attention-based motif learning process, which aids in selecting these key motifs. Once these important motifs are identified, we employ a motif-based graph generator trained to produce explanations based on these motifs. Our experimental results show that our method reliably produces explanations featuring complex molecular structures, achieving complete validity across various molecular datasets. Furthermore, both quantitative and qualitative evaluations demonstrate that the explanation molecules created by our method are more representative and human-understandable than those generated by baseline methods.
2 Preliminary
Notations. A molecular dataset can be depicted as $\mathcal{G}=\{G_{1},G_{2},...,G_{i},...,G_{n}\}$, where each molecular graph within this set is denoted by $G_{i}=(\textbf{V},\textbf{E})$. The label of each graph will be assigned by $l_{i}\in L=\{l_{1},l_{2},...\}$ In this representation, V constitutes a set of atoms, each treated as a node, while E comprises the bonds, represented as edges connecting these nodes. The structural information of each graph $G_{i}$ is encoded in several matrices. The adjacency matrix $\textbf{A}\in\{0,1\}^{N\times N}$ represents the connections between atoms, where a ‘1’ indicates the presence of a bond between a pair of atoms, a ‘0’ signifies no bond, and $N$ represents the total number of atoms in the graph. Additionally, an atom feature matrix $\textbf{X}\in\mathbb{R}^{N\times D}$ captures the features of each atom, with $D$ being the number of features characterizing each atom. Similarly, a bond feature matrix $\textbf{Z}\in\mathbb{R}^{N\times D_{b}}$ is used to describe the attributes of the bonds, where $D_{b}$ denotes the number of features describing each bond.
Graph Neural Networks. Graph Neural Networks (GNNs) leverage adjacency, node feature, and edge feature matrices to learn node features. Despite the existence of various GNN variants like Graph Convolutional Networks (GCNs), Graph Attention Networks (GATs), and Graph Isomorphism Networks (GINs), they generally follow a message-passing mechanism. In this process, each GNN layer aggregates information from a node’s neighbors to refine its representation. For a given hidden layer $i$, the message passing operation can be formulated as $h_{i}=f(h_{i-1},A,Z)$, where $h_{0}=X$. Here, $h_{i}\in\mathbb{R}^{N\times D}$ is the node representation output from layer $i$, and $D$ is the dimension of the output features. The function $f$ involves computing messages based on the previous layer’s node embeddings ($h_{i-1}$), aggregating these messages from neighboring nodes, and updating the hidden representation $h_{i}$ for each node using the aggregated messages.
Motif. A motif is essentially a recurring subgraph within a graph, closely associated with the graph’s property, which has been thoroughly researched across various domains, including biochemistry, ecology, neurobiology, and engineering [24, 31, 1, 2] and has been demonstrated to be of significant importance. GSN [7] introduces motif information into node features to improve the learning of GNNs. HMGNN [43] generates a heterogeneous motif graph to enhance the molecular representation learning. MotifExplainer [44] applies motifs to learn an instance-level explanation for GNNs.
Figure 2: An illustration of the proposed MAGE framework. Given a dataset, a motif extraction algorithm is initially employed to identify all potential motifs. Each motif’s feature encoding is derived from the output encoding produced by the target model, which uses the motif graph as its input. A single-layer attention operator is employed to learn the optimal motif combination, maximizing the likelihood that the target classifier will classify both the reconstructed and original molecular encodings identically. To get a score matrix, the method performs a dot product between the attention coefficient matrix and the prediction probability matrix. This score matrix is then normalized using a degree matrix. Finally, motifs whose corresponding scores exceed a specific threshold are selected.
3 Motif-based Model-Level GNN Explainer
This work introduces a Motif-bAsed GNN Explainer (MAGE), a novel approach for model-level GNN explanations. This method uses motifs as fundamental elements for interpreting the overarching behavior of a trained GNN for molecular graph classification at the model level. Instead of generating a graph at the atomic level, using motifs allows the generator to assemble graphs from these predefined blocks. This method prevents the creation of potentially invalid intermediate substructures and speeds up the generation process. It achieves this by reducing the number of configurations the generator must consider, focusing solely on the arrangement of motifs rather than the specific placement of each atom and bond.
Given a dataset, denoted as $\mathcal{G}$ with $N$ molecules and $C$ classes, our objective is to generate model-level explanations for a target GNN, which comprises a feature extractor $\phi(\cdot)$ and a classifier $f(\cdot)$. $\phi(\cdot)$ takes a molecular graph or subgraph as input and outputs its feature representation. $f(\cdot)$ utilizes this representation to predict the classification of a molecule.
MAGE begins with identifying all potential motifs within $\mathcal{M}$. Following this, an attention-driven motif learning phase is employed to identify the most significant motifs for each class using $\phi(\cdot)$ and $f(\cdot)$. Finally, a graph generator uses the identified motifs to explain each class.
3.1 Motif Extraction
There are mainly three methods for motif extraction. It’s important to note that these motif extraction methods can be integrated into our MAGE explanation method, demonstrating its versatility.
1) Ring and bonded pairs (R&B). Methods [15, 43, 7] in this category recognize simple rings and bonded pairs as motifs. 2) Molecule decomposition. In this category, methods recognize motifs through molecule breakdown. These methods initially establish a catalog of target bonds, either through domain-specific knowledge, such as the “breakable bonds” used by RECAP [21, 50] and BRICS [8, 48, 14], or through hand-crafted rules, like “bridge bonds” [16]. The algorithms remove these target bonds from the molecules and use the consequent molecular fragments as motifs. 3) Molecule tokenization. Certain methods [9, 19] leverage the string representations of molecules, such as SMILES, and directly apply the WordPiece algorithm from the NLP domain. The resulting strings within the constructed vocabulary are regarded as motifs. 4) Data-driven method. MotifPiece [45] statistically identifies underlying structural or functional patterns specific to a given molecular data.
3.2 Class-wise Motif Identification
With extracted significant motifs from the previous section, this section introduces an attention-based motif identification approach for each classification class.
Stage 1: molecule-motif relation score calculation.
Given the absence of a direct linkage between classification categories and motifs, we propose bridging them through molecules. To this end, we employ an attention operator to learn motif-molecule relation scores using the trained feature extractor $\phi(\cdot)$ and a classifier $f(\cdot)$.
For a given molecule graph $G_{i}$ and its associated motif set $M_{i}$, we use an attention operator to identify a combination of motifs that can accurately reconstruct the representations of the molecule graph. The learned attention scores between $G_{i}$ and motifs in $M_{i}$ are interpreted as their relation scores. The initial step involves obtaining feature encoding for the molecule and motif graphs. This is done by feeding a molecule graph or motif graph into $\phi(\cdot)$, and the output encoding serves as the molecule or motif:
$\boldsymbol{h}_{G_{i}}=\phi(G_{i}),\boldsymbol{h}_{m_{i}}=\phi(m_{i}),\text{ for all $m_{i}$}\in M_{i}.$
Next, an attention operator is used to aggregate the motif encoding using $\boldsymbol{h}_{G_{i}}$ as query and $\boldsymbol{h}_{m_{i}}$s as keys and values. The resulting encoding, denoted as $\boldsymbol{h}^{\prime}_{G_{i}}$, is the aggregated molecular encoding by combining motifs in $M_{i}$.
$e_{G_{i}m_{i}}=g(\boldsymbol{W}\boldsymbol{h}_{G_{i}},\boldsymbol{W}\boldsymbol{h}_{m_{i}}),\alpha_{G_{i}m_{i}}=\frac{\exp(e_{G_{i}m_{i}})}{\sum_{m_{i}\in M_{i}}\exp(e_{G_{i}m_{i}})},\boldsymbol{h}^{\prime}_{G_{i}}=\sum_{m_{i}\in M_{i}}\alpha_{G_{i}m_{i}}\cdot\boldsymbol{h}_{m_{i}},$
where $g(\cdot)$ is an attention operator, and $\boldsymbol{W}$ is weight matrix. The target classifier $f(\cdot)$ then uses aggregated encoding to generate predictions. The training process is designed to minimize the cross-entropy loss between the predicted probabilities of the aggregated molecular encoding and those of the original molecular encoding:
$\mathcal{L}=\sum_{i}\mathcal{L}(f(\boldsymbol{h}_{G_{i}}),f(\boldsymbol{h}^{\prime}_{G_{i}})).$ (1)
The attention operator is trained to identify important motifs to represent the molecule. This leads to the molecule-motif relation score $\boldsymbol{\alpha}_{G_{i}m_{i}}$ and the probabilities associated with label predictions $\boldsymbol{\hat{y}}=f(\boldsymbol{h}^{\prime}_{G_{i}})$.
Stage 2: class-motif relation score calculation.
We repeat the process of Stage 1 to compute the molecule-motif scores for each molecular graph. After that, we create a molecule-motif score matrix $\boldsymbol{S}^{mm}\in\mathbb{R}^{V\times N}$ and a molecule-class probability matrix $\boldsymbol{P}\in\mathbb{R}^{N\times C}$. Here, ${S}^{mm}_{pk}$ represents the molecule-motif relation score between molecule $k$ and motif $p$. If the motif $i$ does not appear in the molecule $j$, we set ${S}^{mm}_{ij}=0$. Element ${P}_{kr}$ denotes the probability that molecule $k$ is associated with label $r$.
Subsequently, we perform a matrix multiplication between $\boldsymbol{S}^{mm}$ and $\boldsymbol{P}$ to get class-motif relation score matrix $\boldsymbol{S}^{cm}$:
$\boldsymbol{S}^{cm}=\boldsymbol{S}^{mm}\boldsymbol{P},{S}^{cm}_{pr}=\sum_{k=1}^{N}{S}^{mm}_{pk}{P}_{kr}$
which represents the relationship between motif $p$ and class $r$. Then, we normalize $\boldsymbol{S}^{cm}$ by dividing each column, corresponding to a motif, by the number of molecules it appears.
Stage 3: class-wise motif filtering.
We filter motifs for each class using class-wise motif scores from the previous stage. The motif set for the class $C_{r}$ is
$M_{C_{r}}=\{m_{i}|\boldsymbol{S}^{cm}_{ir}>\theta,1\leq i\leq V\}$ (2)
where $\theta$ is a hyper-parameter to control the size of the important motif vocabulary. An illustration of our class-wise motif identification is provided in Figure 2.
3.3 Class-wise Motif-based Graph Generation
Figure 3: Illustration of class-wise motif-based graph generation.
Section 3.2 identifies motifs of high relevance for each class. Building on this, our methodology employs a motif-based generative model to construct model-level explanations tailored to each class. The motif-based molecular generation method can be divided into tree-constrained [15, 16] and non-constrained methods [41]. In our approach, we employ a tree-constrained method as a generator. However, our approach can easily adapt to non-constrained methods. The graph generation process is illustrated in Figure 3.
3.3.1 Tree Decomposition
Given a target class $r$, a set of important motifs $M_{C_{r}}$ associated with the class $r$, and the set of molecules $\mathcal{G}_{r}$ whose predicted class by trained GNN is $r$. For each molecular graph $G_{i}\in\mathcal{G}_{r}$, we first identify the motifs in $M_{C_{r}}$ from $G_{i}$: $M^{r}_{i}=M_{i}\cap M_{C_{r}}$. Subsequently, we pinpoint every edge in $G_{i}$ that does not belong to any motif in $M^{r}_{i}$. We treat each motif in $M^{r}_{i}$ and each non-motif edge as a cluster node. Then, we construct a cluster graph by connecting clusters with intersections. Finally, a spanning tree $\mathcal{T}$ is selected from this cluster graph, defined as the motif junction tree for $G_{i}$.
3.3.2 Graph Encoder
We encode the latent representation of $G=(\textbf{A},\textbf{X})$ by the target GNN feature extractor $\phi(\cdot)$: $\boldsymbol{h}_{G}=\phi(\textbf{A},\textbf{X})$. The mean $\boldsymbol{\mu}_{G}$ and log variance $\log\boldsymbol{\sigma}_{G}$ of the variational posterior approximation are computed from $\boldsymbol{h}_{G}$ with two separate affine layers. $\boldsymbol{z}_{G}$ is sampled from a Gaussian distribution $\mathcal{N}(\boldsymbol{\mu}_{G},\boldsymbol{\sigma}_{G})$. In the variational posterior approximation, we use two different affine layers to compute the mean, represented as $\boldsymbol{\mu}_{G}$, and the log variance, denoted as $\log\boldsymbol{\sigma}_{G}$, from $\boldsymbol{h}_{G}$. Then, we sample $\boldsymbol{z}_{G}$ from a Gaussian distribution $\mathcal{N}(\boldsymbol{\mu}_{G},\boldsymbol{\sigma}_{G})$.
3.3.3 Tree Encoder
We employ a graph neural network to encode the motif tree $\mathcal{T}=(\boldsymbol{A}_{\mathcal{T}},\boldsymbol{X}_{\mathcal{T}})$, where every node $i$’s feature $x_{i}$ is initially encoded by feeding its corresponding motif graph to $\phi(\cdot)$. Formally, the node embedding and tree embedding are updated as follows:
$\boldsymbol{H}_{\mathcal{T}}=\text{GNN}(\boldsymbol{A}_{\mathcal{T}},\boldsymbol{X}_{\mathcal{T}}),\boldsymbol{h}_{\mathcal{T}}=\text{AGG}(\boldsymbol{H}_{\mathcal{T}}).$
where $\text{GNN}(\cdot)$ is a graph neural network, and $\text{AGG}(\cdot)$ is an aggregation function. $\boldsymbol{z}_{\mathcal{T}}$ is sampled similarly as in the graph encoder.
3.3.4 Tree Decoder
A tree decoder is used to decode a junction tree $\mathcal{T}$ from a tree encoding $\boldsymbol{z}_{\mathcal{T}}$. The decoder builds the tree in a top-down approach, where nodes are created sequentially, one after the other. The node under consideration first obtains its embedding by inputting the existing tree into the tree encoder, and the resulting output is then used as the node embedding. The node embedding is utilized for two predictions: determining whether the node has a child and identifying the child’s label if a child exists. Formally, for node $i$
$\boldsymbol{H}^{\prime}=\text{GNN}(\boldsymbol{A}^{\prime},\boldsymbol{X}^{\prime}),\boldsymbol{p}_{i}=\text{PRED}(\text{COMB}(\boldsymbol{z}_{\mathcal{T}},\boldsymbol{H}^{\prime}_{i})),\boldsymbol{q}_{i}=\text{PRED}^{l}(\text{COMB}^{l}(\boldsymbol{z}_{\mathcal{T}},\boldsymbol{H}^{\prime}_{i}))$
where $\boldsymbol{A}^{\prime}$ and $\boldsymbol{X}^{\prime}$ are the existing tree’s adjacent matrix and node features, respectively. $\text{COMB}(\cdot)$ and $\text{COMB}^{l}(\cdot)$ are combination functions, and $\text{PRED}(\cdot)$ and $\text{PRED}^{l}(\cdot)$ are predictors. When the current node does not have a new child, the decoder will go to the next node in the existing tree.
3.3.5 Graph Decoder
The final step is reproducing the molecular graph from the predicted junction tree $\mathcal{T}$. Let $\mathcal{G}(\mathcal{T})$ be a set of graphs whose junction tree is $\mathcal{T}$. The decoder aims to find $\hat{G}$ as
$\hat{G}=\arg\max_{G^{\prime}\in\mathcal{G(\mathcal{T})}}f^{a}(G^{\prime}).$
In our approach, we aim to ensure that the reproduced molecule maintains the characteristics of the target GNN. So we design the score function as $f^{a}(G_{i})=f(\phi(G_{i}))[r]$, where $r$ is the target label.
3.3.6 Loss Function
The loss function for graph generation comprises two components: the reconstruction loss and the property loss. The reconstruction loss ensures that the generated graph matches the distribution of the original dataset. The property loss ensures that the behavior of the generated graph aligns with the target GNN.
The reconstruction loss can be formally defined as below:
$\mathcal{L_{R}}=\sum\mathcal{L}_{T}+\sum\mathcal{L}_{L}$ (3)
, where $\mathcal{L}_{T}$ corresponds to predicting whether a node has a child, while $\mathcal{L}_{L}$ represents the loss associated with predicting the label of the newly added child. We perform teacher forcing for the reconstruction loss to make the information get the correct history at each step.
The property loss is like below:
$\mathcal{L_{P}}=\sum\text{MSE}(\boldsymbol{h}_{G},\boldsymbol{h}_{\mathcal{T}})+\sum\mathcal{L}(f(\boldsymbol{h}_{\mathcal{T}}),r)$ (4)
, where $\boldsymbol{h}_{\mathcal{T}}$ is the embedding of the generated tree, $\boldsymbol{h}_{G}$ is the corresponding graph embedding, $f(\cdot)$ is the target classifier, and $r$ is the target label. The final loss function is
$\mathcal{L}=\mathcal{L_{R}}+\mathcal{L_{P}}$ (5)
4 Experiment
We conduct qualitative and quantitative experiments on six real-world datasets to evaluate the effectiveness of our proposed methods.
4.1 Dataset and Experimental Setup
We evaluate the proposed methods using molecule classification tasks on six real-world datasets: Mutagenicity, PTC_MR, PTC_MM, PTC_FM, AIDS, and NCI-H23. The details of six datasets and experimental settings are represented in Appendix A and B.
Table 1: Validity results on six real-world datasets. OOM refers to out of memory.
Datasets Models Label 0 Label 1 Average
Mutagenicity XGNN 0.34 0.25 0.295
GNNInterpreter 0.11 0.21 0.16
Ours 1.00 1.00 1.00
PTC_MR XGNN 0.70 0.21 0.45
GNNInterpreter 0.10 0.05 0.07
Ours 1.00 1.00 1.00
PTC_MM XGNN 0.49 0.56 0.51
GNNInterpreter 0.15 0.20 0.17
Ours 1.00 1.00 1.00
PTC_FM XGNN 0.48 0.37 0.42
GNNInterpreter 0.13 0.08 0.10
Ours 1.00 1.00 1.00
AIDS XGNN 0.92 0.71 0.81
GNNInterpreter 0.09 0.08 0.08
Ours 1.00 1.00 1.00
NCI-H23 XGNN OOM OOM OOM
GNNInterpreter 0.59 0.62 0.60
Ours 1.00 1.00 1.00
Table 2: Quantitative results for different explanation methods. We highlight the average class probability (higher is better) for six datasets and the training time for each method on each dataset. The best performances on each dataset are shown in bold. OOM refers to out of memory.
Datasets Models Label 0 Label 1 Average
Mutagenicity XGNN 0.8992 $\pm$ 0.0835 0.9831 $\pm$ 0.0514 0.9411
GNNInterpreter 0.8542 $\pm$ 0.3198 0.9938 $\pm$ 0.0191 0.9240
Variant 0.9897 $\pm$ 0.0754 0.9888 $\pm$ 0.0302 0.9892
Ours 0.9977 $\pm$ 0.0032 0.9941 $\pm$ 0.0240 0.9959
PTC_MR XGNN 0.9906 $\pm$ 0.0718 0.9698 $\pm$ 0.0417 0.9802
GNNInterpreter 0.9067 $\pm$ 0.1728 0.9697 $\pm$ 0.0211 0.9382
Variant 0.9902 $\pm$ 0.0480 0.9641 $\pm$ 0.1163 0.9771
Ours 0.9961 $\pm$ 0.0375 0.9918 $\pm$ 0.0621 0.9939
PTC_MM XGNN 0.9899 $\pm$ 0.0994 0.9266 $\pm$ 0.1059 0.9582
GNNInterpreter 0.9601 $\pm$ 0.0638 0.9541 $\pm$ 0.0207 0.9571
Variant 0.9784 $\pm$ 0.0475 0.9693 $\pm$ 0.0279 0.9738
Ours 0.9914 $\pm$ 0.0342 0.9833 $\pm$ 0.0019 0.9873
PTC_FM XGNN 0.9967 $\pm$ 0.0309 0.9380 $\pm$ 0.0991 0.9673
GNNInterpreter 0.9945 $\pm$ 0.0147 0.9460 $\pm$ 0.0295 0.9702
Variant 0.9882 $\pm$ 0.0024 0.9790 $\pm$ 0.0024 0.9836
Ours 0.9979 $\pm$ 0.0024 0.9890 $\pm$ 0.0024 0.9934
AIDS XGNN 0.9259 $\pm$ 0.1861 0.9977 $\pm$ 0.0225 0.9618
GNNInterpreter 0.4600 $\pm$ 0.4983 0.9973 $\pm$ 0.0116 0.7286
Variant 0.9802 $\pm$ 0.1112 0.9939 $\pm$ 0.0564 0.9870
Ours 0.9883 $\pm$ 0.0663 0.9903 $\pm$ 0.0539 0.9893
NCI-H23 XGNN OOM OOM OOM
GNNInterpreter 0.9883 $\pm$ 0.0711 0.9997 $\pm$ 0.0015 0.9940
Variant 0.9874 $\pm$ 0.0685 0.9863 $\pm$ 0.01069 0.9868
Ours 0.9936 $\pm$ 0.0329 0.9934 $\pm$ 0.0422 0.9935
Baselines. We compare our MAGE model with two state-of-the-art baselines: XGNN and GNNInterpreter. Noted that all methods are tested using official implementation and compared in a fair setting. We extract motifs by identifying “bridge bonds”. We build a variant that uses the same motif extraction and graph generation processes but without our attention-based motif identification step. This variant is to demonstrate the effectiveness of our class-wise motif identification process. The details of this variant will be discussed in section 4.2.
Evaluation metrics. Our quantitative evaluation uses two distinct metrics to assess performance: Validity and Average Probability.
Validity. This metric is defined as the proportion of chemically valid molecules out of the total number of generated molecules [5].
Validity=#valid molecules#generated $\text{Validity}=\frac{\#\,\text{valid molecules}}{\#\,\text{generated molecules}}$
This metric serves as a critical indicator of the practical applicability of our method, ensuring that the generated molecules not only align with the intended class but also adhere to fundamental chemical validity criteria.
Average Probability. Following the GNNInterpreter [37], this metric calculates the average class probability and the standard deviation of these probabilities across 1000 explanation graphs for each class within all six datasets.
Average Probability=class probability#generated $\text{Average Probability}=\frac{\sum\text{class probability}}{\#\,\text{generated molecules}}.$
This approach provides a comprehensive measure of the consistency and reliability of the explanation graphs in representing different classes.
Together, these two metrics offer a robust framework for evaluating the effectiveness of our approach from both the accuracy and applicability perspectives. Considering XGNN’s limitation in generating explanation graphs with edge features, we assign the default bond type from the RDKit library as the default edge feature for explanations produced by XGNN. We also report the training time and sampling time of different models in appendix C.
4.2 Quantitative Results
Table 1 details the Validity results for various models across six datasets. A notable observation from these results is that our method consistently produces valid molecular explanation graphs for both classes in all cases. This reliability stems from our method’s use of valid substructures as foundational elements, effectively overcoming the limitation seen in atom-by-atom generation methods, which often compel the model to create chemically invalid intermediate structures.
Additionally, the data reveals that XGNN exhibits better validity than GNNInterpreter. The reason for this difference lies in the design of the algorithms: XGNN incorporates a manual constraint on the maximum atom degree, which aids in maintaining chemical validity. In contrast, GNNInterpreter is to align the embedding of its generated explanations with the average embedding of the dataset, which does not inherently ensure chemical validity and leads to more invalid generated molecules.
These findings from the validity experiment highlight a principle: maximizing the inductive biases inherent in the molecular graph structure is important. It enables models to achieve higher levels of validity in the outputs, demonstrating that careful consideration of the structural characteristics of molecules can significantly enhance the performance and reliability of these models.
Table 2 shows the average probability of four models on six real-world datasets. Our method outperforms the baseline methods in six out of eight experiments with an average improvement of $1.21\%$, clearly highlighting its effectiveness. Additionally, our method consistently delivers strong results across both classes in each dataset. This contrasts XGNN and GNNInterpreter, which tend to excel in one class but fall short in the other. We attribute this success of our method to its unique strategy of generating explanations in a motif-by-motif way, effectively reducing the influence of extraneous or noise atoms in the data.
Furthermore, we conducted a comparative analysis between our method and a variant model. This variant model operates by first identifying all potential motifs within the dataset. Once these motifs are extracted, it assesses them by putting each motif as an input to the target model. The evaluation of each motif is based on the output probability associated with a particular class, which is assigned as the motif’s score. Following this scoring process, the method ranks all motifs in order of their scores. Only motifs whose scores exceed 0.9 are chosen to refine the selection.
The outcomes of our study indicate that our method outperforms the variant model across all classes, achieving an average enhancement of 2.11%. This significant improvement underscores the efficiency of our attention-based motif learning module. Unlike the variant model, which solely focuses on the individual motifs without considering their functional roles within the molecules, our method learns and understands the interplay between motifs and molecules. This enables our approach to effectively identify and select motifs relevant to the specific molecular context while simultaneously filtering out motifs that do not contribute meaningfully, often called noise motifs.
Table 3: The qualitative results for Mutagenicity dataset. The first row of the table displays the explanation graphs corresponding to the Nonmutagen label, while the second row presents the explanation graphs for the Mutagen label. The final column in the table provides examples of actual graphs. The different colors in the nodes represent different values in the node feature
Class Generated Model-Level Explanation Graphs for Mutagenicity Dataset
XGNN GNNInterpreter Ours Example
Nonmutagen [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Mutagen [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
4.3 Qualitative Results
Table 3 presents the qualitative assessment of three different methods applied to the two classes of the Mutagenicity dataset. The table also includes actual graph examples from the dataset for reference. Our analysis reveals that our method can consistently produce human-understandable and valid explanation graphs, which XGNN and GNNInterpreter seem to lack. More qualitative results can be found in Appendix E.
We observed a trend in GNNInterpreter to produce explanations that are disconnected. This is likely due to its approach of aligning explanation embeddings with the dataset’s average embedding, resulting in a focus on node features rather than structural details. XGNN, on the other hand, tends to create explanations with invalid ring structures. Although XGNN imposes constraints on the degree of each atom, its atom-by-atom generation approach makes it challenging to form valid ring structures. Our method can generate explanations with meaningful substructures like rings.
This inability of XGNN and GNNInterpreter to effectively identify complex molecular structures may contribute to their lower validity in explanation generation. Our method, in contrast, successfully navigates these complexities, offering more meaningful, accurate and structurally coherent explanations. This difference highlights the importance of considering both node features and molecular structures in the generation of explanation graphs, a balance that our approach seems to strike effectively.
5 Related Work
Model-level explanation methods in the field are currently under-researched, with only a few studies addressing this issue. These methods can be divided into two categories. The first category is concept-based methods. GCExplainer [23] introduced the incorporation of concepts into GNN explanations, using the k-Means clustering algorithm on GNN embeddings to identify clusters, each representing a concept. PAGE [32] provides clustering on the embedding space to discover propotypes as explanations. GLGExplainer [3] adopts prototype learning to identify data prototypes and then uses these in an E-LEN model to create a Boolean formula that replicates the GNN’s behavior. GCNeuron [40], on the other hand, employs human-defined rule templates in natural language and considers graphs with certain masked nodes as concepts, using compositional concepts with the highest scores for global explanations. The second category comprises generation-based methods. XGNN [46] uses deep reinforcement learning to generate explanation graphs node by node. GNNInterpreter [37], alternatively, learns a probabilistic generative graph model that identifies the most discriminative graph patterns for explanations.
6 Conclusion
This work proposes a novel motif-based model-level explanation method for Graph Neural Networks, specifically applied to molecular datasets. Initially, the method employs a motif extraction technique to identify all potential motifs. Following this, a one-layer attention mechanism is used to evaluate the relevance of each motif, filtering out those that are not pertinent. Finally, a graph generator is trained to produce explanations based on the chosen motifs. Compared to two baseline methods, XGNN and GNNInterpreter, our proposed method offers explanations that are more human-understandable, accurate, and structurally consistent.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our abstract and introduction accurately introduce the contributions and the experimental results of our paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [N/A]
  9. Justification: Our paper does not have any strong assumptions. We also compare our efficiency with other baselines and show that we have a better efficiency. Our method does not have privacy and fairness issues.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: Our paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide detailed experimental settings in the appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We include our codes in supplemental material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide a detailed experimental setting in the appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report standard deviation for each single experiment.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report the training and sampling time in appendix. We also report the hardware configuration of our experiment in appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have checked the NeurIPS code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Our method is an explanation method on graph dataset, it will not directly have any societal impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our proposed method is an explanation method which has no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We mentioned related works in the reference.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We detailed explained the datasets and baseline models in our paper. We also submit our code as an anonymized zip file.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: our paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: our paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
Appendix A Details of Datasets
Table 4: Statistics and properties of four real-world molecule datasets.
Mutag PTC_MR PTC_MM PTC_FM AIDS NCI-H23
# Nodes (avg) 30.32 14.29 14.32 14.48 15.69 26.07
# Edges (avg) 30.77 14.69 13.97 14.11 16.20 28.10
# Graphs 4337 344 336 349 2000 40353
# Classes 2 2 2 2 2 2
The statistics and properties of four datasets are summarized in Table 4.
Mutagenicity [17, 27] is a chemical compound dataset containing 4,337 molecule graphs. Each of these graphs is classified into either mutagen or non-mutagen categories, signifying their mutagenic impact on the Gram-negative bacterium Salmonella typhimurium.
PTC_MR, PTC_MM, and PTC_FM [25] are three collections of chemical compounds, each respectively documenting carcinogenic effects on male rats (MR), male mice (MM), and female mice (FM).
AIDS [25] comprises 2,000 graphs that represent molecular compounds, all of which are derived from the AIDS Antiviral Screen Database of Active Compounds.
NCI-H23 [25] is a collection of 40353 molecular compounds to evaluate the efficacy against lung cancer cells.
Appendix B Experimental settings
By following baseline approaches, our experiments adopt a simple GCN model and focus on explanation results. For the target GNN, we use a 3-layer GCN as a feature extractor and a 2-layer MLP as a classifier on all datasets. The model is pre-trained and achieves reasonable performances on all datasets. Following the [39], the hidden dimension is selected from $\{16,64\}$. We employ mean-pooling as the readout function and ReLU as the activation function. We use Adam optimizer for training. The target model is trained for 100 epochs, and the learning rate is set to 0.01. The CPU in our setup is an AMD Ryzen Threadripper 2990WX, accompanied by 256 GB of memory, and the GPU is an RTX 4090.
Appendix C Study of training and inference time
Table 5: Training and inference time for different explanation methods. XGNN and GNNInterpreter do not have training time because they only have sampling process.
Datasets Models Training Time Sampling Time
Mutagenicity XGNN 54s
GNNInterpreter 29s
Variant 528s 0.0015s
Ours 539s 0.0029s
PTC_MR XGNN 62s
GNNInterpreter 27s
Variant 338s 0.0005s
Ours 368s 0.0006s
PTC_MM XGNN 46s
GNNInterpreter 17s
Variant 303s 0.0023s
Ours 367s 0.0002s
PTC_FM XGNN 47s
GNNInterpreter 12s
Variant 307s 0.0027s
Ours 386s 0.0018s
AIDS XGNN 19s
GNNInterpreter 32s
Variant 468s 0.0065s
Ours 513s 0.0076s
NCI-H23 XGNN OOM
GNNInterpreter 21s
Variant 10189s 0.0115s
Ours 13657s 0.0126s
We also present a detailed comparison of the training and sampling times between our method and other baseline approaches in table 5. Here, sampling time is the average time to generate a single explanation. Compared to the two baselines, a key distinction of our method lies in its requirement for an initial training phase for the explanation graph generator. This phase confers a significant benefit: the ability to efficiently sample explanations. Our experimental data underscores this advantage. Following the completion of the training phase for the explanation generator, our method demonstrated the capacity to generate a large number of explanations rapidly. Specifically, it could sample 1000 explanations in a matter of seconds. This starkly contrasts other baseline methods, which required at least 3.3 hours to produce the same number of examples. This dramatic reduction in sampling time not only highlights the efficiency of our method but also suggests its practical applicability in scenarios where quick generation of explanations is crucial. Due to the space limit, we put the study of hyper-parameter $\theta$ in Appendix D.
Appendix D Study of Hyperparameter
Table 6: Study of Hyperparameter $\theta$ on Mutagen class of Mutagenicity dataset
$\mathbf{\theta=0.01}$ $\mathbf{\theta=0.05}$ $\mathbf{\theta=0.10}$ $\mathbf{\theta=0.20}$
# Selected Motifs 1591 1328 1206 1073
Average Probability 0.9621 $\pm$ 0.1151 0.9850 $\pm$ 0.0784 0.9977 $\pm$ 0.0032 0.9753 $\pm$ 0.0981
Table 6 shows the study of hyperparameter $\theta$, we found that as $\theta$ increases, the number of motifs selected for the Mutagen class decreases. The average probability rises until $\theta$ reaches 0.10, where peak performance is observed, after which it starts to decline. This trend occurs because when $\theta$ is low, some irrelevant motifs are included in the explanations. As $\theta$ increases, these ’noise’ motifs are filtered out, enhancing performance. However, if $\theta$ is set too high, the resulting small number of selected motifs is insufficient for forming a comprehensive explanation.
Appendix E More Qualitative Results
Table 7: More qualitative results for Nonmutagen class on Mutagenicity dataset.
Class XGNN GNNInterpreter Ours Example
Nonmutagen [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 8: More qualitative results for Mutagen class on Mutagenicity dataset.
Class XGNN GNNInterpreter Ours Example
Mutagen [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]